Question 3. A tire manufacturer warranties its tires to last at least 20,000 miles or “you get a new set of tires.” In its experience, a set of these tires lasts on average 26,000 miles with a standard deviation of 5,000 miles. Assume that the wear is normally distributed. The manufacturer profits $200 on each set sold, and replacing a set costs the manufacturer $400.

(a) What is the probability that a set of tires wears out before 20,000 miles?

(b) What is the probability that the manufacturer turns a profit on selling a set to one customer?

(c) If the manufacturer sells 500 sets of tires, what is the probability that it earns a profit after paying for any replacements? Assume that the purchases are made around the country and that the drivers experience independent amounts of wear.


1
Expert's answer
2021-01-03T14:27:50-0500

Let "X=" the tire life in miles: "X\sim N(\mu, \sigma^2)"

Given μ = 26000 miles, σ = 5000 miles

(a)

"P(X<20000)=P(Z<\dfrac{20000-26000}{5000})"

"=P(Z<-1.2)\approx 0.11507"


The probability that a set of tires wears out before 20,000 miles is "0.11507."


(b)


"P(X\geq20000)=1-P(X<20000)"

"\approx1-0.11507=0.88493"

The probability that the manufacturer turns a profit on selling a set to one customer is "0.88493."


(c) "X\sim N(\mu,\sigma^2/n)"


"P(X\geq20000)=1-P(X<20000)"

"=1-P(Z<\dfrac{20000-26000}{5000/\sqrt{500}})"

"\approx 1-P(Z<-26.8328)\approx1-0=1"

The probability that it earns a profit after paying for any replacements is 1.